OpenStudy (anonymous):

Solve by substitution -x-y-z=-8 -4x+4y+5z=7 2x+2z=4 Hey I just need some one to help me get started on this question. I just need the first couple of steps and then I think I got it. I am more used to doing elimination so I just need a quick refresher to help me get started. Thanks!

OpenStudy (anonymous):

First, you need to get one variable by itself. So use #3. Subtract 2z to get 2x=-2z+4. then divide by 2 to get x by himself. so you get??

OpenStudy (anonymous):

subtract 2z to get 2x =-2z+4...then divide by 2 to get x

OpenStudy (anonymous):

then you get x=-z+2 this becomes another one of your equations. then you plug it in. so then take equation #1 that then becomes ???

OpenStudy (anonymous):

Okay I have x =-2+2 & y=4 so now all I do is plug those into equation 1 2 or 3 and I will have the answer for variable z I think? Thanks you guys are lifesavers!